OpenStudy (anonymous):

You are going to purchase a new car, but being a responsible consumer means doing a little bit of research first. First, you find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3% Using the function A(t) =P(1+r/n)^nt create the function that represents your new car loan that is compounded monthly. The principle will be the price of the vehicle you selected, not how much you are putting down. @ranga

OpenStudy (anonymous):

please help me @ranga

OpenStudy (anonymous):

I already have the equation set up but I need your help for something! @ranga

OpenStudy (anonymous):

The equation I got was 39145(1+0.03/12)^12t

OpenStudy (ranga):

That is correct. Just need to simplify the number within the parenthesis.

OpenStudy (anonymous):

So then it would be be 39145(1.0025)^12T correct?

OpenStudy (ranga):

Yes. \[\Large A=39145(1.0025)^{12t}\]

OpenStudy (anonymous):

The Next question asks Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Set A = 10000 and solve for t.

OpenStudy (ranga):

\[ 10,000 = 39,145(1.0025)^{12t}\\ \frac {10,000}{39,145} = (1.0025)^{12t} \\ 0.2555 = (1.0025)^{12t} \\ \text{Take logarithm on both sides and solve for t} \]

OpenStudy (ranga):

For the previous question where the principal is down to $10,000 don't they tell you how much each monthly payment is?

OpenStudy (anonymous):

I don't think so but when I solved for t i got -45.546. @ranga

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

That answers that question correct?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

I am having trouble though with the last one.

OpenStudy (anonymous):

Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=[(1+r/n^1/c]^cnt where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1+r/n

OpenStudy (anonymous):

So then it would be A(t)=P[1+0.03/12]^1/4(4)(12)t

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

Then i would simplify 1+0.03/12 correct?

OpenStudy (anonymous):

I solved it though and got 1.00000144675926

OpenStudy (anonymous):

That was the answer I got when I solved.

OpenStudy (anonymous):

Is it correct?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Can you type this formula A(t)=[(1+r/n^1/c]^cnt using the Equation Editor so it is clearer?

OpenStudy (ranga):

Or post a screenshot / pic of the formula or the whole problem.

OpenStudy (anonymous):

ok ill try. thanks

OpenStudy (anonymous):

would the answer be for the 3rd question 0.022509321 i didn't round it but this is what i got on the calculator @ranga

OpenStudy (anonymous):

?? @ranga

OpenStudy (anonymous):

can you please help me ranga

OpenStudy (ranga):

I am not sure I understand the third part of the question and hence I am not able to help with that part.

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

@robtobey @Callisto @ganeshie8

OpenStudy (ranga):

My guess is (1+r/n)^(1/c) should be expressed in the form 1 + x/n where x is the adjusted interest rate; r = 0.03, n = 12, c = 4.

OpenStudy (anonymous):

Did you ever find this question out?

OpenStudy (anonymous):

did you ever get the answer to 3? if so could you help me with it?